OpenStudy (anonymous):

Algebra Urgent!!!: Suppose the area of a square is now 400 square meters. If the area is increased by 100%, by what percent will the measure of each edge (side) increase?

OpenStudy (ash2326):

@Maniqueen123 How is the area of a square related to its sides?

OpenStudy (anonymous):

Um idk, but like the area for a square is s^2 so like 400= 20^2 and the formula you use for this question because its percent is x+P(x)=y but m not really sure... :/

OpenStudy (ash2326):

You're right but it's quiet simple What's 100% of 400?

OpenStudy (anonymous):

400

OpenStudy (anonymous):

400*1=400

OpenStudy (anonymous):

so is it like 400+(1)(400)=800

OpenStudy (anonymous):

and than you do like the square root of 800...?

OpenStudy (ash2326):

Yeah so the new area is 400+400=800 find the side now, and then find the % increase in side \[\frac{new\ side-\ old\ side}{old\ side}\times 100 \]

OpenStudy (ash2326):

@Diyadiya help @Maniqueen123 I gotta go

OpenStudy (anonymous):

@Diyadiya, yeah if you're not busy can you please help me im really confused and have an exam tommorow :P

OpenStudy (diyadiya):

(side)^2= 800 side = sqrt(800)

OpenStudy (anonymous):

which equals 20sqrt(2)

OpenStudy (diyadiya):

just a minute let me go through the question again

OpenStudy (anonymous):

kk sure :)

OpenStudy (diyadiya):

Ok Percentage increase = \[\frac{new\ side-\ old\ side}{old\ side}\times 100\] New side = 20sqrt2 Old side =20

OpenStudy (diyadiya):

\[\frac{ 20\sqrt{2} -20}{20} \times 100\]\[\frac{20(\sqrt{2}-1)}{20} \times 100=\frac{\cancel{20}(\sqrt{2}-1)}{\cancel{20}} \times 100= (\sqrt{2}-1) \times 100\]

OpenStudy (diyadiya):

What is the values of \(\sqrt{2} \) ?

OpenStudy (anonymous):

umm its like 1.alot of numbers

OpenStudy (diyadiya):

Did you understand till here?

OpenStudy (anonymous):

i think for the answer you can leave it like tht

OpenStudy (anonymous):

yea :) So i think you can write it like p= (sqrt(2)-1)%

OpenStudy (diyadiya):

\(\sqrt{2}=1.414\) \((1.414 -1) \times 100 =0.414 \times 100 =41.4 \) % Is that right? @ash2326

OpenStudy (ash2326):

Yeah it's correct!! Great work @Diyadiya

OpenStudy (diyadiya):

Thanks :)

OpenStudy (diyadiya):

@Maniqueen123 understood?

OpenStudy (anonymous):

yea thanks alot guys :) you really helped me <3

OpenStudy (diyadiya):

You're Welcome :)